# Boulevard du Triomphe, B-1050 Brussels, Belgium

### v1

@@ Abstract
High-Frequency Gravitational Waves (HFGWs) constitute a unique window on the early Universe as well as exotic astrophysical objects. If the current gravitational wave experiments are more dedicated to the low frequency regime, the graviton conversion into photons in a strong magnetic field constitutes a powerful tool to probe HFGWs. In this paper, we show that neutron stars, due to their extreme magnetic field, are a perfect laboratory to study the conversion of HFGWs into photons. Using realistic models for the galactic neutron star population, we calculate for the first time the expected photon flux induced by the conversion of an isotropic stochastic gravitational wave background in the magnetosphere of the ensemble of neutron stars present in the Milky Way. We compare this photon flux to the observed one from several telescopes and derive upper limits on the stochastic gravitational wave background in the frequency range $10^{8}\,\rm Hz$ $10^{25}\,\rm Hz$. We find our limits to be competitive in the frequency range $10^{8}\,\rm Hz$ $10^{15}\,\rm Hz$.
@@ 1 Introduction
The search for low-frequency (nHz-kHz) Gravitational Waves (GWs) has been strongly motivated by theoretical predictions in the last decades. This has naturally led to a series of experimental projects aiming to detect GWs below the $\sim\rm kHz$. Among them, the Laser Interferometer Gravitational-Wave Observatory (LIGO) LIGOScientific:2016aoc first observed GWs in the kHz range from binary black holes. More recently, the Pulsar Timing Array (PTA) community (NANOGrav NANOGrav:2023gor , PPTA Reardon:2023gzh , EPTA EPTA:2023fyk , and IPTA Antoniadis:2022pcn ) has claimed strong evidence of a Stochastic Gravitational Wave Background (SGWB) in the nHz regime. The latter is strongly motivated by supermassive black holes binaries Middleton:2020asl ; Phinney:2001di . Early Universe sources have been shown to offer a similar level of explanation to the SGWB Vagnozzi:2020gtf ; Chen:2021ncc ; Sakharov:2021dim ; Benetti:2021uea ; Ashoorioon:2022raz ; Wu:2023pbt ; IPTA:2023ero ; Wu:2023dnp ; Dandoy:2023jot ; Madge:2023dxc .
@@
In analogy to the electromagnetic spectrum that covers orders of magnitudes in frequency, GWs could be produced over a large frequency range, each of them being associated with different physical phenomena. Interestingly, GWs with frequencies larger than the ones currently searched for in experiments are difficult to motivate with known astrophysical objects or standard cosmology Aggarwal:2020olq . Any detection of such GWs would, therefore, be a hint for exotic early Universe sources; phase transitions, preheating after inflation, oscillons, cosmic strings, or exotic astrophysical objects, for instance, primordial black holes Ghiglieri:2015nfa ; Wang:2019kaf ; Ghiglieri:2020mhm ; Ghiglieri:2022rfp ; Khlebnikov:1997di ; Easther:2006gt ; Garcia-Bellido:2007nns ; Anantua:2008am ; Dolgov:2011cq ; Dong:2015yjs (see Ref. Caprini:2018mtu for a review of different sources). These primordial GWs are expected to propagate freely in the Universe to form a SGWB, as in the case of the background potentially observed by the PTA.
@@
However, High-Frequency Gravitational Waves (HFGWs) have been significantly less explored than the low-frequency range. The main reason for this relies on the fact that GW detectors are usually sensitive to GWs with wavelengths comparable to their size. Targeting HFGWs, therefore, requires a radically different experimental setup (see Refs. Nishizawa:2007tn ; Grishchuk:2003un ; Braginsky:1978yp ; Aggarwal:2020olq ; Ito:2019wcb ; Ito:2020wxi for a non-exhaustive list of recent proposals).
@@
Nonetheless, the detection of HFGWs may be possible via the (inverse) Gertsenshtein effect Raffelt:1987im ; Boccaletti:1970pxw : similarly to axions, GWs in the presence of a magnetic field can convert into photons. This has naturally led to use the existing axion experiments as HFGW detectors and translating the existing constraints on the axion to photon coupling into constraints on the GWs abundance $\Omega_{\rm gw}$ (usually expressed in terms of the characteristic strain $h_{c}$ingwald:2020ist ; Ejlli:2019bqj ; Berlin:2021txa ; Domcke:2022rgu ; Tobar:2022pie .
@@
One of the main reason that makes astrophysical environments particularly suitable for observing the Gertsenshtein effect lies in the natural presence of strong magnetic fields (which enhance the graviton-photon conversion) impossible to produce in the laboratory. Among the most extreme examples, neutron stars (NSs) are known to posses a surface magnetic field around $B\approx 10^{13}\,\rm Gauss$ Faucher-Giguere:2005dxp ; Popov:2009jn and, therefore, appear to be one of the most efficient places to convert gravitons into photons. Furthermore, NS magnetospheres are relatively well understood and allow for reliable predictions. In the presence of a SWGB the ensemble of NSs present in the Milky Way—estimated to be around $10^{9}$ NSs Ofek:2009wt ; Sartore:2009wn —would then induce a photon flux at the same frequency as the SGWB, as illustrated in Fig. 1. The galactic NS population therefore behaves as a powerful tool to probe the presence of a high frequency SGWB.
Motivated by the last statement, we calculate in this work the radiated photon flux from the conversion of a high frequency SGWB in the magnetosphere of the ensemble of neutron stars (NSs) within the Milky Way (see Ref. Safdi:2018oeu for a similar study in the axion case). Using recent models for the NS magnetic field, spin period, and spatial distribution Faucher-Giguere:2005dxp ; Popov:2009jn , we compare this photon flux to the observations conducted by several telescopes (see Ref. Hill:2018trh for a summary). Finally we extract competitive constraints on the amount of GWs in a large frequency range, from $\sim 10^{8}\,\rm Hz$ to $\sim 10^{25}\,\rm Hz$.
In the following, we first start in Sec. 2 by recalling the inverse Gertsenshtein effect in the specific context of neutron star magnetospheres. We describe in Sec. 3 the expected induced photon flux from the galactic NSs and detail the method used to extract constraints on the GW abundance. Finally, we show our results in Sec. 3.1 and conclude in Sec. 4. The specific models for the NS distributions are discussed in App. A.
@@ 2 Graviton-Photon Conversion in Neutron Stars
Each term describes a different physical effect: $\Delta_{\mathrm{pla},\parallel}=-\frac{m_{\mathrm{pla}}^{2}}{2\omega}$ and $\Delta_{\mathrm{pla},\perp}=-\frac{m_{\rm{pla}}^{2}\omega^{2}}{2\omega(\omega^{2}-\omega_{c})}$ are the plasma corrections to the photon mass, with $\omega_{c}\approx 10^{19}\left(\frac{B}{10^{12}{\rm G}}\right)$ Raffelt:1987im . The difference between the two polarizations comes from the Cotton-Mouton effect Ejlli:2018hke ; Ejlli:2018ucq ; Ito:2023fcr . $\Delta_{\mathrm{vac},\|(\perp)}=7(4)\frac{\alpha\omega}{90\pi}\left(\frac{B_{t}}{B_{c}}\right)^{2}$ is the QED vacuum correction. We defined $m_{\mathrm{pla}}^{2}=\frac{4\pi\alpha n_{c}}{m_{c}}$, $B_{c}=\frac{m_{e}^{2}}{e}$ and $\omega=2\pi f$, with $f$ the graviton/photon frequency. Moreover, $\alpha$ and $e$ are the QED fine structure constant and the electron charge. Finally, $n_{c}$ and $m_{c}$ are the number density and invariant mass of the particles in the plasma. In NSs we have $n_{c}\approx n_{e}$ and $m_{c}\approx m_{e}$, the electron number density and the electron mass, respectively.
@@
with $f$ the frequency of the graviton/photon and considering a generic GW path in the magnetic field between $\ell_{0}$ and $\ell_{1}$ [^1] In what follows, we assume that the typical wavelenght of the GW is much smaller than the NS magnetosphere. This, in addition to the large occupation number, allows us to describe GWs as made of individual gravitons with defined trajectories. A more precise study of the problem using a wave desciption of the system is left for future works..
In the specific case of NSs, the magnetic field extends from its surface $r_{0}$ to the end of the magnetosphere $R\approx 2\pi/T$ Ito:2023fcr , with $T$ the rotation period. Assuming equipartition and averaging over the directions of the magnetic field we parametrize the component of the magnetic field perpendicular to the GW trajectory in the NS magnetosphere as Ito:2023fcr
@@ 2.1 Radial Trajectory
$P_{0,\|}^{\mathrm{high-f}}\approx\frac{2(\Delta_{\rm M}(r_{0})r_{0})^{2}}{5^{6/5}(\Delta_{\rm{vac,\parallel}}(r_{0})r_{0})^{4/5}}\left|\Gamma\left(\frac{2}{5}\right)-\Gamma\left(\frac{2}{5},-\frac{i\Delta_{\rm{vac,\parallel}}(r_{0})r_{0}}{5}\right)\right|^{2},$ (20)
@@
$f_{\mathrm{tra}}\approx 3.59\times\left(\frac{m_{e}}{e}\right)^{3/2}\sqrt{\frac{2\pi}{B_{0}T}}.$ (22)
For frequencies around $f_{\mathrm{tra}}$, it exists a region on the graviton path, not too far from the NS surface, where $\Delta_{\mathrm{pla},\parallel}$ and $\Delta_{\mathrm{vac},\parallel}$ compensate so that the photon effective mass vanishes. When this mass is small enough, the graviton-photon conversion satisfies both momentum and energy conservation and is enhanced. The maximum of the conversion probability would therefore be around $f_{\mathrm{tra}}$. A good approximation of the maximum conversion probability is given by
@@
Figure 2: Conversion probability as a function of the frequency for an NS with $B_{0}=10^{8}$ Gauss, $T=0.1$ s, and $r_{0}=10$ km (left) and for $B_{0}=10^{13}$ Gauss, $T=1$ s, and $r_{0}=10$ km. The curves respectively in dark and light green are the analytical approximation at low frequencies and high frequencies for the parallel polarization. The yellow curve is the analytical approximation for the perpendicular polarization and the dashed gray line is the maximum conversion probability approximation for the parallel polarization. Finally, the black curve in the right panel is the numerical result for the parallel polarization.
@@
For the $\perp$ polarization, the conversion probability is simply given by Eq. (20) (replacing $\Delta_{\mathrm{vac},\parallel}$ by $\Delta_{\mathrm{vac},\perp}$). Since $\Delta_{\mathrm{vac},\perp}\sim f$, the effective photon becomes smaller as we lower the frequency. Therefore, the conversion probability increases as the frequency gets lower, until it reaches a maximum when the effective mass becomes negligible.
@@ 2.2 Generic Trajectory
Figure 3: Left: Conversion probability for the perpendicular polarization as a function of the impact parameter. Green and yellow lines show the probability for $f=10^{12}\,\rm Hz$ and $f=10^{15}\,\rm Hz$, respectively. Vertical and horizontal dashed lines show the value of $b_{c}(f)$ and $P_{0}^{\rm plateau}$ for each frequency, respectively. The magnetic field is set to $B_{0}=10^{13}\,\rm G$. Right: Critical impact parameter as a function of the frequency. Green and yellow lines show $B_{0}=10^{12}\,\rm G$ and $B_{0}=10^{13}\,\rm G$, respectively. Solid and dashed lines show the perpendicular and parallel polarizations, respectively. In both plots, the NS spin period and radius has been taken to be $T=1$ s, and $r_{0}=10$ km.
@@ 3 Photon flux from GW conversion in galactic Neutron Stars
where $M_{\rm pl}$ is the reduced Planck mass. We have introduced the characteristic strain $h_{c}$ Maggiore:2000gv ,
@@
From NS observations, different models for their evolution and properties have been developed. We choose two of them Faucher-Giguere:2005dxp ; Popov:2009jn , which differ in the evolution of the magnetic field but are consistent with data (see for instance Refs. Manchester:2001fp ; Edwards:2001de ). While Faucher-Giguere:2005dxp has a constant magnetic field, Popov:2009jn accounts for its decay and predicts a weaker magnetosphere for dead NS. Then, each model predicts a different distribution for both variables. Independently of the model, the probability distribution for both variables inside of each subgroup is well represented by a log-normal distribution, and so we distinguish $P_{\rm act.}(B_{0},T)$ for active NSs from $P_{\rm dead}(B_{0},T)$ for dead NSs. It is estimated that 0.4% of the galactic NS population are active, while 99.6% are dead Safdi:2018oeu . Further details on the spatial and probability distributions can be found in App. A.
@@
Figure 4: Left: Predicted frequency spectrum of the total induced flux $\partial F^{\rm galac.}_{\gamma}/\partial f$, normalized by $1/h_{c}^{2}$. Blue and green lines show the results for a NS model with constant and decaying $\vec{B}$, respectively. We set $r_{0}=10\,\mathrm{km}$. Right: Frequency spectrum measured on Earth, adapted from Ref. Hill:2018trh . The thick band corresponds to the 95% confidence interval.

### v2

@@ Abstract
High-Frequency Gravitational Waves (HFGWs) constitute a unique window on the early Universe as well as exotic astrophysical objects. While the current gravitational wave experiments are more dedicated to the low frequency regime, the graviton conversion into photons in a strong magnetic field constitutes a powerful tool to probe HFGWs. In this paper, we show that neutron stars, due to their extreme magnetic field, are a perfect laboratory to study the conversion of HFGWs into photons. Using realistic models for the galactic neutron star population, we calculate for the first time the expected photon flux induced by the conversion of an isotropic stochastic gravitational wave background in the magnetosphere of the ensemble of neutron stars present in the Milky Way. We compare this photon flux to the observed one from several telescopes and derive upper limits on the stochastic gravitational wave background in the frequency range $10^{8}\,\rm Hz$ $10^{25}\,\rm Hz$. We find our limits to be competitive in the frequency range $10^{8}\,\rm Hz$ $10^{12}\,\rm Hz$.
@@ 1 Introduction
The search for low-frequency (nHz-kHz) Gravitational Waves (GWs) has been strongly motivated by theoretical predictions in the last decades. This has naturally led to a series of experimental projects aiming to detect GWs below the $\sim\rm kHz$. Among them, the Laser Interferometer Gravitational-Wave Observatory (LIGO) LIGOScientific:2016aoc first observed GWs in the kHz range from binary black holes. More recently, the Pulsar Timing Array (PTA) community (NANOGrav NANOGrav:2023gor , PPTA Reardon:2023gzh , EPTA EPTA:2023fyk , and IPTA Antoniadis:2022pcn ) has claimed strong evidence of a Stochastic Gravitational Wave Background (SGWB) in the nHz regime. The latter is strongly motivated by supermassive black holes binaries Middleton:2020asl ; Phinney:2001di . Early Universe sources have been shown to offer a similar level of explanation to the SGWB Vagnozzi:2020gtf ; Chen:2021ncc ; Sakharov:2021dim ; Benetti:2021uea ; Ashoorioon:2022raz ; Wu:2023pbt ; IPTA:2023ero ; Wu:2023dnp ; Dandoy:2023jot ; Madge:2023dxc ; Kawai:2023nqs ; Vagnozzi:2023lwo .
@@
In analogy to the electromagnetic spectrum that covers orders of magnitudes in frequency, GWs could be produced over a large frequency range, each of them being associated with different physical phenomena. Interestingly, GWs with frequencies larger than the ones currently searched for in experiments are difficult to motivate with known astrophysical objects or standard cosmology Aggarwal:2020olq . Any detection of such GWs would, therefore, be a hint for exotic early Universe sources; phase transitions, preheating after inflation, oscillons, cosmic strings, or exotic astrophysical objects, for instance, primordial black holes Ghiglieri:2015nfa ; Wang:2019kaf ; Ghiglieri:2020mhm ; Ghiglieri:2022rfp ; Khlebnikov:1997di ; Easther:2006gt ; Garcia-Bellido:2007nns ; Anantua:2008am ; Dolgov:2011cq ; Dong:2015yjs ; Costa:2022lpy ; Costa:2022oaa ; Gehrman:2023esa (see Ref. Caprini:2018mtu for a review of different sources). These primordial GWs are expected to propagate freely in the Universe to form a SGWB, as in the case of the background potentially observed by the PTA.
@@
However, High-Frequency Gravitational Waves (HFGWs) have been significantly less explored than the low-frequency range. The main reason for this relies on the fact that GW detectors are usually sensitive to GWs with wavelengths comparable to their size. Targeting HFGWs, therefore, requires a radically different experimental setup (see Refs. Nishizawa:2007tn ; Grishchuk:2003un ; Braginsky:1978yp ; Aggarwal:2020olq ; Ito:2019wcb ; Ito:2020wxi ; Kahn:2023mrj for a non-exhaustive list of recent proposals).
@@
Nonetheless, the detection of HFGWs may be possible via the (inverse) Gertsenshtein effect Raffelt:1987im ; Boccaletti:1970pxw : similarly to axions, GWs in the presence of a magnetic field can convert into photons. This has naturally led to use the existing axion experiments as HFGW detectors and translating the existing constraints on the axion to photon coupling into constraints on the GWs abundance $\Omega_{\rm gw}$ (usually expressed in terms of the characteristic strain $h_{c}$ingwald:2020ist ; Ejlli:2019bqj ; Berlin:2021txa ; Domcke:2022rgu ; Tobar:2022pie , see Fig. 6.
@@
One of the main reasons that makes astrophysical environments particularly suitable for observing the Gertsenshtein effect lies in the natural presence of strong magnetic fields (which enhance the graviton-photon conversion) impossible to produce in the laboratory. Among the most extreme examples, neutron stars (NSs) are known to possess a surface magnetic field around $B\approx 10^{13}\,\rm Gauss$ Faucher-Giguere:2005dxp ; Popov:2009jn and, therefore, appear to be one of the most efficient places to convert gravitons into photons. Furthermore, NS magnetospheres are relatively well understood and allow for reliable predictions. In the presence of a SWGB the ensemble of NSs present in the Milky Way—estimated to be around $10^{9}$ NSs Ofek:2009wt ; Sartore:2009wn —would then induce a photon flux at the same frequency as the SGWB, as illustrated in Fig. 1. The galactic NS population therefore behaves as a powerful tool to probe the presence of a high frequency SGWB.
Motivated by the last statement, we calculate in this work the radiated photon flux from the conversion of a high frequency SGWB in the magnetosphere of the ensemble of NSs within the Milky Way (see Ref. Safdi:2018oeu for a similar study in the axion case). Using recent models for the NS magnetic field, spin period, and spatial distribution Faucher-Giguere:2005dxp ; Popov:2009jn , we compare this photon flux to the observations conducted by several telescopes (see Ref. Hill:2018trh for a summary). Finally we extract competitive constraints on the amount of GWs in a large frequency range, from $\sim 10^{8}\,\rm Hz$ to $\sim 10^{25}\,\rm Hz$.
In the following, we first start in Sec. 2 by recalling the inverse Gertsenshtein effect in the specific context of NS magnetospheres. We describe in Sec. 3 the expected induced photon flux from the galactic NSs and detail the method used to extract constraints on the GW abundance. Finally, we show our results in Sec. 3.1 and conclude in Sec. 4. The specific models for the NS distributions are discussed in App. A.
@@ 2 Graviton-Photon Conversion in Neutron Stars
Each term describes a different physical effect: $\Delta_{\mathrm{pla},\parallel}=-\frac{m_{\mathrm{pla}}^{2}}{2\omega}$ and $\Delta_{\mathrm{pla},\perp}=-\frac{m_{\rm{pla}}^{2}\omega^{2}}{2\omega(\omega^{2}-\omega_{c}^{2})}$ are the plasma corrections to the photon mass, with $\omega_{c}\approx 10^{19}\left(\frac{B}{10^{12}{\rm G}}\right)\,\mathrm{Hz}$ Raffelt:1987im . The difference between the two polarizations comes from the Cotton-Mouton effect Ejlli:2018hke ; Ejlli:2018ucq ; Ito:2023fcr . $\Delta_{\mathrm{vac},\|(\perp)}=7(4)\frac{\alpha\omega}{90\pi}\left(\frac{B_{t}}{B_{c}}\right)^{2}$ is the QED vacuum correction. We defined $m_{\mathrm{pla}}^{2}=\frac{4\pi\alpha n_{c}}{m_{c}}$, $B_{c}=\frac{m_{e}^{2}}{e}$ and $\omega=2\pi f$, with $f$ the graviton/photon frequency. Moreover, $\alpha$ and $e$ are the QED fine structure constant and the electron charge. Finally, $n_{c}$ and $m_{c}$ are the number density and invariant mass of the particles in the plasma. In NSs we have $n_{c}\approx n_{e}$ and $m_{c}\approx m_{e}$, the electron number density and the electron mass, respectively.
@@
considering a generic GW path in the magnetic field between $\ell_{0}$ and $\ell_{1}$ [^1] In what follows, we assume that the typical wavelength of the GW is much smaller than the NS magnetosphere. This, in addition to the large occupation number, allows us to describe GWs as made of individual gravitons with defined trajectories. A more precise study of the problem using a wave description of the system is left for future works..
In the specific case of NSs, the magnetic field extends from its surface $r_{0}$ to the end of the magnetosphere $R\approx T/(2\pi)$ Ito:2023fcr , with $T$ the rotation period. Assuming equipartition and averaging over the directions of the magnetic field we parametrize the component of the magnetic field perpendicular to the GW trajectory in the NS magnetosphere as Ito:2023fcr
@@ 2.1 Radial Trajectory
$P_{0,\|}^{\mathrm{high-f}}\approx\frac{2(\Delta_{\rm M}(r_{0})r_{0})^{2}}{5^{6/5}(\Delta_{\rm{vac,\parallel}}(r_{0})r_{0})^{4/5}}\left|\Gamma\left(\frac{2}{5}\right)-\Gamma\left(\frac{2}{5},-\frac{i\Delta_{\rm{vac,\parallel}}(r_{0})r_{0}}{5}\right)\right|^{2},$
@@
$f_{\mathrm{tra}}\approx 6.5\times\frac{m_{e}^{3/2}}{e}\sqrt{\frac{2\pi}{B_{0}T}}\,.$ (21)
For frequencies around $f_{\mathrm{tra}}$, there is a region where $\Delta_{\mathrm{pla},\parallel}$ and $\Delta_{\mathrm{vac},\parallel}$ compensate, and thus the photon effective mass vanishes. This region is close to the NS surface. When this mass is small enough, the graviton-photon conversion satisfies both momentum and energy conservation and is enhanced. The maximum of the conversion probability would therefore be around $f_{\mathrm{tra}}$. A good approximation of the maximum conversion probability is given by
@@
Figure 2: Conversion probability as a function of the frequency for an NS with $B_{0}=10^{8}$ Gauss, $T=0.1$ s, and $r_{0}=10$ km (left) and for $B_{0}=10^{13}$ Gauss, $T=1$ s, and $r_{0}=10$ km (right). The curves respectively in dark and light green are the analytical approximation at low frequencies and high frequencies for the parallel polarization. The yellow curve is the analytical approximation for the perpendicular polarization which matches the numerical result. The dashed gray line is the maximum conversion probability approximation for the parallel polarization and the black dashed line is $f_{tra}$. Finally, the black curve in the right panel is the numerical result for the parallel polarization.
@@
For the $\perp$ polarization, the conversion probability is simply given by Eq. (2.1) (replacing $\Delta_{\mathrm{vac},\parallel}$ by $\Delta_{\mathrm{vac},\perp}$). Since $\Delta_{\mathrm{vac},\perp}\sim f$, the effective photon becomes smaller as we lower the frequency. Therefore, the conversion probability increases as the frequency gets lower, until it reaches a maximum when the effective mass becomes negligible. The numerical results match the analytical approximation.
@@ 2.2 Generic Trajectory
Figure 3: Left: Conversion probability for the perpendicular polarization as a function of the impact parameter. Green and yellow lines show the probability for $f=10^{9}\,\rm Hz$ and $f=10^{12}\,\rm Hz$, respectively. Vertical and horizontal dashed lines show the value of $b_{c}(f)$ and $P_{0}^{\rm plateau}$ for each frequency, respectively. The magnetic field is set to $B_{0}=10^{13}\,\rm G$. Right: Critical impact parameter as a function of the frequency. Green and yellow lines show $B_{0}=10^{12}\,\rm G$ and $B_{0}=10^{13}\,\rm G$, respectively. Solid and dashed lines show the perpendicular and parallel polarizations, respectively. In both plots, the NS spin period and radius has been taken to be $T=1$ s, and $r_{0}=10$ km.
@@ 3 Photon Flux from GW Conversion in Galactic Neutron Stars
where $M_{\rm pl}$ is the reduced Planck mass, and we have integrated over half a sphere. We have introduced the characteristic strain $h_{c}$ Maggiore:2000gv ,
@@
From NS observations, different models for their evolution and properties have been developed. We choose two of them Faucher-Giguere:2005dxp ; Popov:2009jn , which differ in the evolution of the magnetic field. While Faucher-Giguere:2005dxp has a constant magnetic field, Popov:2009jn accounts for its decay and predicts a weaker magnetosphere for dead NS. Although both models are consistent with data (see for instance Refs. Manchester:2001fp ; Edwards:2001de ), the former leads to stronger constraints and the latter to more conservative ones. We keep both in order to illustrate the sensitivity of our predictions to the models, which is the dominant uncertainty of our results. Then, each model predicts a different distribution for both variables. Independently of the model, the probability distribution for both variables inside of each subgroup is well represented by a log-normal distribution, and so we distinguish $P_{\rm act.}(B_{0},T)$ for active NSs from $P_{\rm dead}(B_{0},T)$ for dead NSs. It is estimated that 0.4% of the galactic NS population are active, while 99.6% are dead Safdi:2018oeu . Further details on the spatial and probability distributions can be found in App. A.
@@
Figure 4: Left: Predicted frequency spectrum of the total induced flux $\partial F^{\rm galac.}_{\gamma}/\partial f$, normalized by $1/h_{c}^{2}$. Blue and green lines show the results for a NS model with constant and decaying $\vec{B}$, respectively. We set $r_{0}=10\,\mathrm{km}$. Right: Frequency spectrum measured on Earth, adapted from Ref. Hill:2018trh . The thick band corresponds to the 95% confidence interval. Data allows to put a bound on $h_{c}$ at each frequency.